Solve by completing the square:
x^4-5x^2-36=0

Answer by checkley75(3666)

X^4-5X^2-36=0
X^4-5X^2+6.25=36+6.25
(X^2-2.5)(X^2-2.5)=42.25
(X^2-2.5)^2=42.25
X^2-2.5=SQRT42.25
X^2-2.5=6.5
X^2-2.5-6.5=0
(X^2-9)=0
(X-3)(X+3)=0
X-3=0
X=3 ANSWER.
X+3=0
X=-3 ANSWER.